# Large Model Empowered Streaming Semantic Communications for Speech Translation

### v1

@@ Abstract
Semantic communications have been explored to perform downstream intelligent tasks by extracting and transmitting essential information. In this paper, we introduce a large model-empowered streaming semantic communication system for speech translation across various languages, named LaSC-ST. Specifically, we devise an edge-device collaborative semantic communication architecture by offloading the intricate semantic extraction module to edge servers, thereby reducing the computational burden on local devices. To support multilingual speech translation, pre-trained large speech models are utilized to learn unified semantic features from speech in different languages, breaking the constraint of a single input language and enhancing the practicality of the LaSC-ST. Moreover, the input speech is sequentially streamed into the developed system as short speech segments, which enables low transmission latency without the degradation in speech translation quality. A novel dynamic speech segmentation algorithm is proposed to further minimize the transmission latency by adaptively adjusting the duration of speech segments. According to simulation results, the LaSC-ST provides more accurate speech translation and achieves streaming transmission with lower latency compared to existing non-streaming semantic communication systems.
@@
Large model, semantic communications, speech translation, streaming speech transmission.
@@ I Introduction
Semantic communications have undergone unprecedented advancements over the past few years due to the booming of artificial intelligence (AI). To contend with the explosive growth of data traffic, deep learning (DL)-enabled semantic communications have been considered a promising solution to provide intelligent data transmission and address numerous bottlenecks in conventional communications [1].
@@
Shannon and Weaver categorized communication into three levels, including syntax communications, semantic communications, and pragmatic communications [2]. The conventional communication paradigm falls under syntax communications, which quantifies information at the bit level and aims to achieve a low bit-error rate (BER) or symbol-error rate (SER). This bit-oriented communication framework ignores the meaning behind the transmission data, running counter to the ultimate goal of semantic exchange in wireless communications. In this context, many researchers have investigated semantic information from different theoretical perspectives [3, 4]. However, a rigorous mathematical formula for quantifying semantics has not yet been established. For the sake of efficient semantic representation, DL techniques have shown their potential to extract semantic information inherent in the source by leveraging the superior learning and fitting capabilities of sophisticated neural networks (NNs). DL-enabled semantic communications have attracted substantial attention and overcome the challenge of approximate semantic representation.
@@
DL-enabled semantic communications explore two transmission goals: source reconstruction and task execution. Particularly, Xie et al. [5] established the pioneering semantic commutation system for text transmission, DeepSC, by jointly designing the semantic and channel coding. Jiang et al. [6] devised a hybrid automatic repeat request (HARQ)-based semantic communication system to strengthen the transmission reliability of semantic information. Inspired by the flow of intelligence, Dong et al. [7] carried out a semantic communication system for image restoration, which enhances model flexibility across diverse transmission scenarios. Additionally, Xie et al. [8] developed task-oriented semantic communications for visual question-answering by fusing textual and visual semantic features at the receiver to infer the context, and Huang et al. [9] introduced an adaptive semantic coding approach to sever object detection task in semantic communications. In [10], Zhang et al. built a unified semantic communication framework for multitask execution by invoking a lightweight feature selection network.
@@
In semantic communications for speech transmission, Weng et al. [11] proposed the first semantic communication system for speech reconstruction, named DeepSC-S. Xiao et al. [12] designed a variant of DeepSC-S to augment the intelligibility of the produced speech by optimizing the rate-distortion trade-off. In [13, 14], Weng et al. further studied intelligent speech tasks in semantic communications, including speech recognition and speech-to-text translation. Although task-oriented semantic communications for speech transmission have demonstrated superiority in serving AI tasks compared to conventional speech transmission protocols, we are still facing the following challenges:
@@ 1. C1:
    The computational resources of user devices are insufficient for implementing complicated feature extraction.
@@ 2. C2:
    Existing works only support speech in one language and lack the adaptability for fine-tuning to other languages.
@@ 3. C3:
This paper proposes a novel large model-empowered semantic communication system for multilingual speech translation, LaSC-ST, to address these challenges, facilitating seamless speech communication across linguistic boundaries and reducing transmission latency by processing the input speech in a streaming manner. The main contributions of this paper can be summarized as follows:
@@
  • Simulation results demonstrate the superiority of LaSC-ST in serving the speech translation task and the effectiveness of the dynamic speech segmentation algorithm.
The rest of this paper is structured as follows. In Section II, the model of the large model-empowered streaming semantic communications is provided. Section III presents the details of the proposed LaSC-ST. Section IV presents experimental results and Section V concludes this paper.
@@ II-A Edge-Device Collaborative Communication Framework
The motivation of this work is to support the real-time speech translation task in semantic communications when users have various linguistic backgrounds. The model structure of the designed system is shown in Fig. 1. From the figure, the system input consists of speech in one of the supported languages, $\boldsymbol{s}=\left[{\boldsymbol{s}}_{1},\;{\boldsymbol{s}}_{2},\;...,\;{\boldsymbol{s}}_{T}\right]$, where ${\boldsymbol{s}}_{t}$ is $t$-th short speech segment in $\boldsymbol{s}$. ${\boldsymbol{s}}_{t}$ is fed into the speech compressor to obtain the intermediate representation, $\boldsymbol{p}$, through a lightweight NN. It is worth mentioning that the following speech segment, ${\boldsymbol{s}}_{t+1}$, is being captured while the speech compressor is processing ${\boldsymbol{s}}_{t}$, and the data size of $\boldsymbol{p}$ is significantly smaller than that of ${\boldsymbol{s}}_{t}$ to streamline the interaction between the local device and the edge server, which addresses the preceding challenge C3. Denote the speech compressor as ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$, then $\boldsymbol{p}$ is written as
@@
The intermediate features, $\boldsymbol{p}$, are transmitted to the pre-trained large speech model on the edge server to extract semantic features $\boldsymbol{f}$ within a short period due to the exceptional computational prowess of edge server. $\boldsymbol{p}$ is downloaded to the local device and mapped to the symbols, $\boldsymbol{x}$, by the channel encoder. $\boldsymbol{x}$ can be presented as
@@
After transmitting the encoded symbols, $\boldsymbol{x}$, over the wireless fading channel, the received symbols, $\boldsymbol{y}$, can be expressed as
@@
The channel decoder at the receiver takes $\boldsymbol{y}$ as input and attains the estimated semantic features, $\widehat{\boldsymbol{f}}$, denoted as
@@
The recovered $\widehat{\boldsymbol{f}}$ is uploaded to the edge server to infer the translated semantic information, $\widehat{\boldsymbol{p}}$, by the large speech model. Then, $\widehat{\boldsymbol{p}}$ is retrieved by the local speech predictor to generate the translated speech segment, $\widehat{\widetilde{\boldsymbol{s}}}_{t}$, expressed as
@@
The translated speech segments are continuously provided to the receiver user to ensure a seamless speech communication experience. It is noteworthy that the speech compressor, channel, encoder, channel decoder, and speech predictor are designed as lightweight NNs to alleviate the computational burden on the local device, which resolves the challenge C1. The state-of-the-art large speech model is an ideal solution to address the challenge C2 because it is capable of extracting coherent semantic features and generating translated outputs across numerous languages. Additionally, the speech compressor and the speech predictor are pre-trained along with the large speech model without accounting for any communication issues, while the channel encoder and the channel decoder are trained under specific channel conditions. The mean-squared error (MSE) is considered as the loss function to minimize the error between $\boldsymbol{f}$ and $\widehat{\boldsymbol{f}}$, modelled as
@@ III Large Model Empowered Semantic Communications for Speech Translation
In this section, we introduce the details of LaSC-ST. Additionally, the proposed dynamic speech segmentation algorithm is presented.
III-A LaSC-ST
The proposed LaSC-ST is shown in Fig. 2. From the figure, the one-dimensional (1D) convolutional neural network (CNN)-based speech compressor condenses a batch of speech segments, ${\boldsymbol{S}}_{t}$, into the preliminary features, $\boldsymbol{P}$. The cutting-edge large speech model, Meta Seamless Communication [16], is deployed on the edge server and returns the learned semantic features, $\boldsymbol{F}$, to the local device. Meta Seamless Communication supports many AI tasks, such as speech recognition and speech translation, across over 100 spoken languages. The dense layer-based channel encoder converts $\boldsymbol{F}$ into symbols $\boldsymbol{X}$ before transmission over the wireless channel.
Figure 2: Model structure of LaSC-ST for end-to-end streaming speech translation.
At the receiver, the obtained symbols, $\boldsymbol{Y}$, are passed through the dense layer-based channel decoder, and the output is the recovered semantic features, $\widehat{\boldsymbol{F}}$. The MSE loss is calculated after the channel decoder and backpropagated to the transmitter to update the trainable NN parameters of the channel encoder and decoder. The training algorithm is described in Algorithm 1. Next, the translated speech segments, $\widehat{\widetilde{\boldsymbol{S}}}_{t}$, are attained by feeding $\widehat{\boldsymbol{F}}$ into the 1D CNN and dense layer-based speech predictor. Finally, $\widehat{\widetilde{\boldsymbol{S}}}_{t}$ at each transmission is continuously concatenated to form the batch of translated speech, $\widehat{\boldsymbol{S}}$.
@@ III-A LaSC-ST
1:Input: Batch of input speech $\boldsymbol{S}$, pre-trained speech compressor ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$ and larger speech model, fading channel $\boldsymbol{H}$, Gaussian noise $\boldsymbol{N}$.
@@
5:         Upload $\boldsymbol{P}$ to the edge server and extract $\boldsymbol{F}$.
6:         Download $\boldsymbol{F}$ to the local device.
@@
10:         Compute ${\mathcal{L}}_{\mathrm{MSE}}(\boldsymbol{\beta},\;\boldsymbol{\gamma})$ to update ${\mathfrak{T}}_{\mathrm{CE}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{CD}}(\cdot)$.
11:     end for
12:end while
13:Output: Trained ${\mathfrak{T}}_{\mathrm{CE}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{CD}}(\cdot)$.
@@ III-B Dynamic Speech Segmentation Algorithm
In the LaSC-ST framework, the input speech is divided into multiple speech segments, each maintaining a constant duration. However, this rigid segmentation approach is not suitable in some extreme scenarios. For instance, when a speech segment contains considerable semantic information, the processes of generating $\boldsymbol{F}$ from $\boldsymbol{P}$ and obtaining $\widehat{\boldsymbol{P}}$ form $\widehat{\boldsymbol{F}}$ become highly time-consuming, thereby increasing the overall communication latency. To combat this, a dynamic speech segmentation algorithm is proposed to enable a more intelligent and adaptive mechanism for partitioning the input speech, which adjusts the current speech segment duration according to the semantic content of the previous segment. It is intuitive that the subsequent speech segment contains essential information if the amplitude of the speech samples in the current segment increases abruptly. Conversely, a decrease in amplitude suggests that the following segment may represent a pause in the speech flow. Additionally, when the amplitude of the current segment approaches zero, it implies that this segment carries no information.
@@
Inspired by this, an example of the dynamic speech segmentation algorithm is shown in Fig. 3. From the figure, the duration of the following segment is longer when the slope of the speech amplitude in the current segment exceeds zero, and it is shorter when the slope is smaller than zero. The silent segment is not fed into the LaSC-ST for inference. Denote the duration of ${\boldsymbol{s}}_{t}$ is $d_{{\boldsymbol{s}}_{t}}$, then $d_{{\boldsymbol{s}}_{t+1}}$ can be expressed as
$d_{{\boldsymbol{s}}_{t+1}}=\left\{\begin{array}[]{l}\max\left[m,\;\left(1-pe^{k}\right)d_{{\boldsymbol{s}}_{t}}\right],\;\;\;\;\;\;\;\;\;\;\;\mathrm{if}\;k>0,\\ \min\left[n,\;\left(1-q\ln\left(k+1\right)\right)d_{{\boldsymbol{s}}_{t}}\right],\;\;\mathrm{else}.\end{array}\right.$ (8)
where $k$ indicates the slope of the speech amplitude in the current segment. $p$ and $q$ are two hyperparameters. $m$ and $n$ denote two thresholds that regulate the duration of speech segments, preventing them from becoming too long or too short, respectively.
According to the dynamic speech segmentation, the testing algorithm of LaSC-ST for enabling streaming speech translation in semantic communications is illustrated in Algorithm 2.

### v2

# Large Model Empowered Streaming Speech Semantic Communications

@@ Abstract
In this paper, we introduce a large model-empowered streaming semantic communication system for speech transmission across various languages, named LSSC-ST. Specifically, we devise an edge-device collaborative semantic communication architecture by offloading the intricate semantic extraction and channel coding modules to edge servers, thereby reducing the computational burden on local devices. To support multilingual speech transmission, pre-trained large speech models are utilized to learn unified semantic features from speech in different languages, breaking the constraint of a single input language and enhancing the practicality of the LSSC-ST. Moreover, the input speech is sequentially streamed into the developed system as short speech segments, which enables low transmission latency without degrading the quality of the produced speech. A novel dynamic speech segmentation algorithm is proposed to further reduce the transmission latency by adaptively adjusting the duration of speech segments. According to simulation results, the LSSC-ST provides more accurate speech transmission and achieves a streaming manner with lower latency compared to the existing non-streaming semantic communication systems.
@@
Large model, semantic communications, streaming speech transmission.
@@ I Introduction
Semantic communications have been proved to undergo unprecedented advancements over the past few years due to the booming of artificial intelligence (AI). To contend with the explosive growth of data traffic, deep learning (DL)-enabled semantic communications have been considered a promising solution to provide intelligent data transmission and address numerous bottlenecks in conventional communications [1, 2].
@@
According to Shannon and Weaver [3], communication can be categorized into three levels, including syntax communications, semantic communications, and pragmatic communications. The conventional communication paradigm falls under syntax communications, quantifies information at the bit level, and aims to achieve a low bit-error rate (BER) or symbol-error rate (SER). This bit-oriented communication framework ignores the meaning behind the transmission data, running counter to the ultimate goal of semantic exchange in wireless communications. In this context, semantic information has been investigated from different theoretical perspectives [4, 5]. For the sake of efficient semantic representation, DL techniques have shown their potential to extract semantic information inherent in the source by leveraging the superior learning and fitting capabilities of sophisticated neural networks (NNs). DL-enabled semantic communications have attracted substantial attention and overcome the challenge on approximate semantic representation [6].
@@
DL-enabled semantic communications explore two transmission goals: source reconstruction and task execution. Particularly, Xie et al. [7] pioneered text semantic commutation system, DeepSC, by jointly designing the semantic and channel coding. Jiang et al. [8] devised a hybrid automatic repeat request (HARQ)-based semantic communication system to strengthen the transmission reliability of semantic information. Inspired by the flow of intelligence, Dong et al. [9] carried out a semantic communication system for image restoration, which enhances model flexibility across diverse transmission scenarios. Additionally, Xie et al. [10] developed task-oriented semantic communications for visual question-answering by fusing textual and visual semantic features at the receiver to infer the context. In [11], Zhang et al. built a unified semantic communication framework for multitask execution by invoking a lightweight feature selection network.
@@
In semantic communications for speech transmission, Weng et al. [12] proposed the first semantic communication system for speech reconstruction, named DeepSC-S. In [13], Weng et al. further studied intelligent speech tasks in semantic communications, such as speech recognition and speech-to-text translation. Although task-oriented semantic communications for speech transmission have demonstrated superiority in serving AI tasks compared to the conventional speech transmission protocols, we are still facing the following challenges:
@@ 1. C1:
    The computational resources of user devices are insufficient for complicated feature extraction.
@@ 2. C2:
    Existing works only support speech in one language and lack the adaptability for fine-tuning to others.
@@ 3. C3:
Large models have been applied to address various challenges in wireless communications [14]. This paper proposes a novel task-oriented large model-empowered semantic communication system for speech transmission, LSSC-ST, to address these challenges. It considers multilingual speech translation tasks, facilitating seamless speech communication across linguistic boundaries and reducing transmission latency by processing the input speech in a streaming manner. The main contributions of this paper can be summarized as follows:
@@
The rest of this paper is structured as follows. In Section II, the model of the large model-empowered streaming semantic communications is provided. Section III presents the details of the proposed LSSC-ST. Section IV presents experimental results and Section V concludes this paper.
@@ II-A Edge-Device Collaborative Communication Framework
The motivation of this work is to support the real-time speech translation task in semantic communications when users have various linguistic backgrounds. The model structure of the designed system is shown in Fig. 1. From the figure, the system input consists of speech in one of the supported languages, $\boldsymbol{s}=\left[{\boldsymbol{s}}_{1},\;{\boldsymbol{s}}_{2},\;...,\;{\boldsymbol{s}}_{T}\right]$, where ${\boldsymbol{s}}_{t}$ is $t$-th short speech segment in $\boldsymbol{s}$. ${\boldsymbol{s}}_{t}$ is fed into the speech compressor on the local device in a streaming manner to obtain the intermediate representation, $\boldsymbol{p}$, through a lightweight NN. It is worth mentioning that the following speech segment, ${\boldsymbol{s}}_{t+1}$, is being captured while the speech compressor is processing ${\boldsymbol{s}}_{t}$, and the data size of $\boldsymbol{p}$ is significantly smaller than that of ${\boldsymbol{s}}_{t}$ to streamline the interaction between the local device and the edge server, which addresses the preceding challenge, C3. Denote the speech compressor as ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$, then $\boldsymbol{p}$ is written as
@@
The intermediate features, $\boldsymbol{p}$, are transmitted to the pre-trained large speech model on the edge server to extract semantic features $\boldsymbol{f}$ within a short period due to the exceptional computational prowess of the edge server. $\boldsymbol{p}$ is mapped to the symbols, $\boldsymbol{x}$, by the channel encoder. $\boldsymbol{x}$ can be expressed as
@@
The encoded symbols, $\boldsymbol{x}$, on the edge server are transmitted through the wireless fading channel. The received symbols, $\boldsymbol{y}$, on another edge server can be expressed as
@@
The channel decoder takes $\boldsymbol{y}$ as input and attains the estimated semantic features, $\widehat{\boldsymbol{f}}$, denoted as
@@
Recovered $\widehat{\boldsymbol{f}}$ on the edge server is converted into the translated semantic information, $\widehat{\boldsymbol{p}}$, by the large speech model. Then, $\widehat{\boldsymbol{p}}$ is downloaded to the local device and retrieved by the speech predictor to generate the translated speech segment, $\widehat{\widetilde{\boldsymbol{s}}}_{t}$, expressed as
@@
The translated speech segments are continuously provided to the receiver user to ensure seamless speech communication. It is noteworthy that the speech compressor and speech predictor are designed as lightweight NNs to alleviate the computational burden on the local device, which resolves challenge C1. The state-of-the-art large speech model is an ideal solution to address challenge C2 because it is capable of extracting coherent semantic features and generating translated outputs across numerous languages. Additionally, the speech compressor and the speech predictor are pre-trained along with the large speech model without accounting for any communication issues while the channel encoder and the channel decoder are trained under specific channel conditions. The mean-squared error (MSE) is considered as the loss function to minimize the error between $\boldsymbol{f}$ and $\widehat{\boldsymbol{f}}$, modelled as
@@ III Large Model Empowered Semantic Communications for Speech Translation
In this section, we introduce the LSSC-ST. Additionally, the proposed dynamic speech segmentation algorithm is presented.
III-A LSSC-ST
The proposed LSSC-ST is shown in Fig. 2. From the figure, the one-dimensional (1D) convolutional neural network (CNN)-based speech compressor condenses a batch of speech segments, ${\boldsymbol{S}}_{t}$, into the preliminary features, $\boldsymbol{P}$. The cutting-edge large speech model, Meta Seamless Communication [16], is deployed on the edge server and returns the learned semantic features, $\boldsymbol{F}$. Meta Seamless Communication supports many AI tasks, such as speech recognition and speech translation, across over 100 spoken languages. The dense layer-based channel encoder converts $\boldsymbol{F}$ into symbols $\boldsymbol{X}$ on the edge server before transmission over the wireless channel.
Figure 2: Model structure of LSSC-ST for end-to-end streaming speech translation.
At the receiver, the obtained symbols, $\boldsymbol{Y}$, are passed through the dense layer-based channel decoder, and the output is the recovered semantic features, $\widehat{\boldsymbol{F}}$. The MSE loss is calculated after the channel decoder and backpropagated to the transmitter to update the trainable NN parameters of the channel encoder and decoder. The training algorithm is described in Algorithm 1. Next, the translated speech segments, $\widehat{\widetilde{\boldsymbol{S}}}_{t}$, are attained by feeding $\widehat{\boldsymbol{P}}$ into the 1D CNN and dense layer-based speech predictor. Finally, $\widehat{\widetilde{\boldsymbol{S}}}_{t}$ at each transmission is continuously concatenated to form a batch of translated speech, $\widehat{\boldsymbol{S}}$.
@@ III-A LSSC-ST
1:Input: Batch of input speech $\boldsymbol{S}$, pre-trained speech compressor ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$ and large speech model, fading channel $\boldsymbol{H}$, Gaussian noise $\boldsymbol{N}$.
@@
5:         Upload $\boldsymbol{P}$ to the edge server.
6:         Extract $\boldsymbol{F}$ from $\boldsymbol{P}$.
@@
10:         Compute ${\mathcal{L}}_{\mathrm{MSE}}(\boldsymbol{\beta},\;\boldsymbol{\gamma})$.
11:         Update $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$.
12:     end for
13:end while
14:Output: Trained ${\mathfrak{T}}_{\mathrm{CE}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{CD}}(\cdot)$.
@@ III-B Dynamic Speech Segmentation Algorithm
In the LSSC-ST framework, the input speech is divided into multiple speech segments, each maintaining a constant duration. However, this rigid segmentation approach is not suitable in some extreme scenarios. For instance, when a speech segment contains considerable semantic information, the processes of generating $\boldsymbol{F}$ from $\boldsymbol{P}$ and obtaining $\widehat{\boldsymbol{P}}$ form $\widehat{\boldsymbol{F}}$ become highly time-consuming, thereby increasing the overall communication latency. To combat this, a dynamic speech segmentation algorithm is proposed to enable a more intelligent and adaptive mechanism for partitioning the input speech, which adjusts the current speech segment duration according to the semantic content of the previous segment. It is intuitive that the subsequent speech segment contains essential information if the amplitude of the speech samples in the current segment increases abruptly. Conversely, a decrease in amplitude suggests that the following segment may represent a pause in the speech flow. Additionally, when the amplitude of the current segment approaches zero, it implies that this segment carries no information.
@@
Inspired by this, an example of the dynamic speech segmentation algorithm is shown in Fig. 3. From the figure, the duration of the following segment is longer when the slope of the speech amplitude in the current segment exceeds zero, and it is shorter when the slope is smaller than zero. Silent segments are excluded from processing by the LSSC-ST during inference. Denote the duration of ${\boldsymbol{s}}_{t}$ to be $d_{{\boldsymbol{s}}_{t}}$, then $d_{{\boldsymbol{s}}_{t+1}}$ can be expressed as
$d_{{\boldsymbol{s}}_{t+1}}=\left\{\begin{array}[]{l}\max\left[m,\;\left(1-pe^{k}\right)d_{{\boldsymbol{s}}_{t}}\right],\;\;\;\;\;\;\;\;\;\;\;\mathrm{if}\;k>0,\\ \min\left[n,\;\left(1-q\ln\left(k+1\right)\right)d_{{\boldsymbol{s}}_{t}}\right],\;\;\mathrm{else},\end{array}\right.$ (8)
where $k$ indicates the slope of the speech amplitude in the current segment, $p$ and $q$ are two hyperparameters, and $m$ and $n$ denote two thresholds that regulate the duration of speech segments, preventing them from becoming too long or too short, respectively.
TABLE I: Parameter settings of the proposed LSSC-ST.
Layer Name Parameters Activation
Speech Compressor 2$\times$1D CNN 128, 64 channels ReLU
Channel Encoder 3$\times$×Dense 3$\times$2048 units None
Channel Decoder 3$\times$×Dense 3$\times$2048 units ReLU
Speech Predictor 2$\times$1D CNN 2$\times$1280 channels ReLU
1$\times$×Dense 1 unit None
Large Speech Model Meta Seamless Communication
According to the dynamic speech segmentation, the testing algorithm of LSSC-ST for enabling streaming speech translation in semantic communications is illustrated in Algorithm 2.